# Cristian Sestito  Shady Agwa  and Themis Prodromakis

### v1

@@ VI Conclusion
In this work, we present a hardware architecture to showcase the TrIM dataflow. The architecture is organized in a hierarchical manner, to effectively cope with convolutional layers. Reconfigurable shift register buffers are included to make the engine agnostic to different ifmap sizes. As a case study, the VGG-16 CNN is considered and an architecture of 1512 processing elements is implemented onto the XCZU7EV FPGA. The achieved peak throughput is 453.6 Giga Operations per Second. When compared to Eyeriss, a state-of-the-art accelerator based on the row stationary dataflow, the TrIM-based architecture guarantees 5.1$\sim$ 5.1$\times$ 5.1 × less memory accesses, confirming the effectiveness of the triangular input movement to maximize data utilization. In addition, the proposed architectures results up to 12.2$\sim$ 12.2$\times$ 12.2 × more energy-efficient than other FPGA counterparts.
@@
  • In the direction of high energy efficiency, we aim to design and tape-out a TrIM-based architecture on ASIC, thus unlocking its applicability in large-scale systems.

### v2

@@ VI Conclusion
In this work, we present a hardware architecture to showcase the TrIM dataflow. The architecture is organized in a hierarchical manner, to effectively cope with convolutional layers. Reconfigurable shift register buffers are included to make the engine agnostic to different ifmap sizes. As a case study, the VGG-16 CNN is considered and an architecture of 1512 processing elements is implemented onto the XCZU7EV FPGA. The achieved peak throughput is 453.6 Giga Operations per Second. When compared to Eyeriss, a state-of-the-art accelerator based on the row stationary dataflow, the TrIM-based architecture guarantees 3$\sim$ 3$\times$ 3 × less memory accesses, confirming the effectiveness of the triangular input movement to maximize data utilization. In addition, the proposed architectures results up to 11.9$\sim$ 11.9$\times$ 11.9 × more energy-efficient than other FPGA counterparts.
@@
  • Insertion of a global buffer for ifmaps and weights, in order to reduce the count of off-chip memory access, thus strengthening the results with respect to previous art.
  • In the direction of high energy efficiency, we are currently working on a tape-out of the TrIM-based architecture using a commercial CMOS technology, thus unlocking the applicability of TrIM in large-scale systems.